The value of Ksp for TlCl(s) in equilibrium with water at 25°C is 1.9 × 10−4 M2. Write the chemical equation that represents the solubility equilibrium for TlCl(s) and calculate its solubility in grams per liter
in water at 25°C.
Answer #1

TlCl(s) -------------> Tl^+(aq) + Cl^-(aq)

---------------------- x -------------- x

Ksp   = [Ti^+][Cl^-]

1.9*10^-4 = x*x

x^2    = 1.9*10^-4

x    = 0.013784

solubility of TlCl   = 0.013784mole/L

molar mass of TlCl   = 239.82g/mole

solubility of TlCl   = 0.013784mole/L * 239.82g/mole

                         = 3.3056g/L >>>>answer